OpenStudy (anonymous):

The nth term of a sequence can be described by the expression three n minus two . What is the fiftieth term? Answer forty-eight fifty fifty-five one hundred forty-eight

OpenStudy (anonymous):

\[n^{th}~term=3n-2\]\[15^{th}~term=3(15)-2\]\[15^{th}~term=30-2\]\[15^{th}~term=~?\]

OpenStudy (lgbasallote):

in my land...fifty usually means 50....

OpenStudy (lgbasallote):

substitute 50 into the n of 3n - 2 so just do 3(50) - 2 and you'll get your answer got it?

OpenStudy (anonymous):

oh sorry i read that wrong :\

OpenStudy (lgbasallote):

haha happens to the best of us =))

OpenStudy (anonymous):

I wish I could give you both medals..

OpenStudy (anonymous):

It's fine. :)

OpenStudy (anonymous):

Which expression results from applying the Identity Property of Multiplication to five u plus five v plus eight v plus v plus ten to get an equivalent expression? Answer five u plus five v plus nine v plus ten five u plus v times the quantity five plus eight end quantity plus v plus ten five times the quantity u plus v end quantity plus eight v plus v plus ten five u plus thirteen v plus v plus ten

OpenStudy (lgbasallote):

hint: identity property of multiplication is times 1 does that help?